OpenStudy (anonymous):

Indefinite Integrals..... indefinite integral (x+1)(3x-2) (dx) 1st thing I did was "distribute" using FOIL. and then I got stuck because of what I have left 3x^2 + x -2 (dx) What next?

jimthompson5910 (jim_thompson5910):

you can break up the integral like this \[\Large \int (3x^2 + x - 2)dx = \int (3x^2)dx+\int (x)dx + \int (-2)dx\]

OpenStudy (anonymous):

@jim_thompson5910 ok that looks good, but what about the 3x^2 part?

jimthompson5910 (jim_thompson5910):

you can pull out the coefficient \[\Large \int (3x^2)dx = 3*\int (x^2)dx\]

OpenStudy (anonymous):

for x(dx) = x^2/2 and for (-2) =-2x

jimthompson5910 (jim_thompson5910):

A lot of rules that you find with differentiating are similar in integration

jimthompson5910 (jim_thompson5910):

correct on both of those

OpenStudy (anonymous):

so everytime I see a number infront of the variable I pull it outside the integral sign as you did above?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\[\Large \int k*f(x)dx = k*\int f(x) dx\] where k is a constant

OpenStudy (anonymous):

this is what I got as my final answer: \[\frac{ 3x^{3} }{ 3 } + \frac{x ^{2} }{ 2 } -2x\]

jimthompson5910 (jim_thompson5910):

the 3x^3 over 3 reduces to x^3

jimthompson5910 (jim_thompson5910):

and don't forget the +C

OpenStudy (anonymous):

oh yes okay, does first one simplify to x^3 ?

jimthompson5910 (jim_thompson5910):

yes it does

OpenStudy (anonymous):

Ok great! Thank you so much!! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome